OpenStudy (anonymous):

Is sin(x)+cos(x)=1 an identity or not an identity?

OpenStudy (apoorvk):

Not an identity, as it is not valid for every x. sin^2(x) + cos^2(x) = 1 IS an identity, however.

OpenStudy (anonymous):

The identity is \[\sin^2 x +\cos^2 x =1\]

OpenStudy (anonymous):

How can I prove that sin(x)+cos(x)=1 is not an identity?

OpenStudy (amistre64):

find a value for x that makes it fail

OpenStudy (amistre64):

x=pi/4 would do fine